OpenStudy (anonymous):

What is the missing integer in this problem? (Do not show work or skip spaces when answering this question. Make sure to type in only the missing integer) -11 + ____ = -9

OpenStudy (anonymous):

can u show us your essay ?

OpenStudy (anonymous):

wat do u mean by essay

OpenStudy (anonymous):

i got -2

OpenStudy (anonymous):

No , -2 doesn't work ! Because : \[-11+(-2)=-13\]

OpenStudy (anonymous):

im confused how to work it

OpenStudy (anonymous):

First, it is clear to you how is -11+(-2)=-13 ?

OpenStudy (anonymous):

yea u add

OpenStudy (anonymous):

ok, now for the qustion, we are looking for a number such that : \[-11 + ...=-9\] we remark that our number is : \[-9-(-11)\]

OpenStudy (anonymous):

a positive 2

OpenStudy (anonymous):

yes ! \[-11+ 2=-9\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

you could have just set it up like this : -11 + x = - 9 (and just solve for x) x = - 9 + 11 x = 2